# ImDy: Human Inverse Dynamics from Imitated Observations

### v1

@@ 1 Introduction
The rapid progress in human motion capture based on computer vision has made an enormous amount of human motion data available to the research community (Mahmood et al., 2019; Mandery et al., 2016). The accumulation of human motion manages to push motion understanding forward in various tasks, including behavior understanding (Punnakkal et al., 2021; Shahroudy et al., 2016) and character animation (Guo et al., 2022; Tevet et al., 2023). However, given the vision-based nature, most current efforts focus only on visible kinematics information. The invisible factors, especially the dynamic factors, which could carry deeper insights into the underlying production mechanism of human motion, are typically overlooked, such as driven torques and ground reaction forces. This limits the current motion understanding algorithms from wider applications to domains where physical constraints must be seriously considered, such as robotics, healthcare, and sports training. To alleviate this, we focus on identifying the driven torques and ground reaction forces for human motion from pure kinematics MoCap data, known as human inverse dynamics (ID).
Human inverse dynamics, as a basic step toward physical motion modeling, has been extensively discussed by the biomechanics community for applications like gait analysis. A fundamental obstacle is that it could not be measured non-intrusively. Therefore, computationally expensive optimization-based methods are widely adopted and mature software is developed (Delp et al., 2007; Damsgaard et al., 2006; Werling et al., 2021). However, accurately measured ground reaction forces are required to ensure a determinate solution, which could be expensive and applicable only in restricted laboratory settings. Also, the optimization process could be sensitive to small disturbances in either motion capture noises or subject variances. These make it hard to scale up for wider applications to general motion. Given the success achieved by data-driven methods in computer vision and natural language processing, deep-learning-based methods are proposed (Zell & Rosenhahn, 2015; Zell et al., 2017; Lv et al., 2016), aiming at scalable human inverse dynamics with only kinematic observations as inputs.
@@
Given this, we project our sights on the recent progress of Imitation Learning (IL) (Luo et al., 2021; 2023), which replicates recorded human motion through fully simulated humanoids with physical control signals, namely, joint torques. A key insight is that with the goal of kinematics phenomenon imitation, IL might also implicitly imitate the dynamics production mechanism, known as ID. However, IL is not directly applicable to ID. Despite the visual resemblances between the recorded and simulated motion, kinematic errors still exist. These errors could be neglected for kinematic analyses, however, when it comes to dynamic analysis, they could be amplified drastically (Uchida & Seth, 2022). Moreover, existing successful IL algorithms are typically based on joint-actuated SMPL (Loper et al., 2015) avatars, whose physical properties and topology differ from real humans. To this end, extracting ID knowledge from IL becomes critical. Here, we adopt the state-of-the-art motion IL algorithm (Luo et al., 2023) and physics simulator (Makoviychuk et al., 2021) to imitate recorded motions, extracting the observed kinematic states, joint torques, and the ground reaction forces, resulting in a large-scale human inverse dynamics database named Imitated Dynamics (ImDy) with more than 150-hour human motion. There are two major merits of ImDy. First, it is scalable. Multiple samples could be concurrently collected in the simulator without expensive laboratory setups, which extends the border of ID data acquisition. As shown in Fig. 1, we could even pair some rather complex motions with ID data, which is hard to achieve in laboratories. Second, it is holistic. Beyond the ground reaction force and ID typically recorded in laboratories for previous efforts (Zell et al., 2020; Mourot et al., 2022; Han et al., 2023), the physics simulator enables us to access the GRFs and joint torques of all human body segments, as shown in Fig. 1.
@@
We demonstrate the efficacy of ImDyS through a wide span of experiments. First, we evaluate our method on ImDy for a basic performance illustration with simulated ImDy. Then ImDyS is evaluated on GroundLink (Han et al., 2023), which contains real-world ground reaction force. Furthermore, we demonstrate the efficacy of ImDy on the recent real-world human dynamics dataset AddBiomechanics (Werling et al., 2024).
@@ 2 Background
Conventional Inverse Dynamics. Inverse dynamics, known as inferring forces/moments from kinematic observations, have been discussed for long in the biomechanics community. In this literature, it is formulated as an optimization problem: given a representative model of a subject, the joint kinematics over time w.r.t. the subject model, and the external forces imposed on the subject, find the driving torques that produce the observed motion (Uchida & Delp, 2021). The Newtonian dynamic equations are involved as
@@ 3.2 Imitated Data Acquisition
AddBiomechanics (Werling et al., 2024) 273 57.6 GRF* & Torques Rajagopal et al. (2016) Real
@@
We adopt AMASS (Mahmood et al., 2019) and KIT (Mandery et al., 2016) as two major data sources. Sequences involving humans interacting with objects other than the ground are excluded, resulting in over 50 hours of motion. All the sequences are re-sampled to 30FPS, with the z-axis as the gravity axis. Then, the sequences are imitated three times by the two primitive policies and the multiplicative policy with a simulation frequency of 60Hz, resulting in over 150 hours of human motion data with dynamics. States including $q,\dot{q},\beta$ are recorded in synchronous with the torque $\tau$, all restored in the format of SMPL (Loper et al., 2015) if possible. Moreover, GRFs for the whole body are also recorded, resulting in ImDy, a large-scale human motion dynamics dataset.
@@ 4.1 Formulation
Motion States $s$ could be represented by either SMPL parameters, joint angles, joint coordinates, or marker coordinates. However, due to the topology divergence, the conversion among SMPL parameters, joint angles, and joint coordinates is non-trivial with limited performances. To guarantee that ImDyS could be seamlessly adopted to both ImDy and real-world biomechanics data, we adopt marker coordinates as motion state representation for ImDyS. The state $s^{t}=(m^{t},\dot{m}^{t})$ is composed of marker coordinates $m^{t}$ and finite-differentiated velocities $\dot{m}^{t}$ at timestamp $t$, which are easy to obtain for both ImDy and AddBiomechanics (Werling et al., 2024). Two temporal windows before and after the transition with a length of $w$ are included for contextual information. Notice that human physical properties like height and weight could also be implicitly represented by the markers. The states are canonicalized w.r.t. the heading direction of $s^{t}$.
@@
Ground Reaction Forces $\lambda$. Different from previous efforts (Mourot et al., 2022; Han et al., 2023; Werling et al., 2024) with foot GRFs only, we predict full-body GRF $\lambda\in\mathbb{R}^{J\times 3}$ as in Fig. 1.
@@ 5.4 Evaluation on AddBiomechanics
Metrics. AddBiomechanics (Werling et al., 2024) is recently proposed with over 50 hours of human dynamics data from 273 subjects. We adopt the armless part of this dataset. We follow the train/test split in Addbiomechanics and report mPJE for the joint torque normalized by body weight.
@@ Appendix B AddBiomechanics Results Analysis
Figure 12: AddBiomechanics data distribution. Among all activities, walking and running account for over 75%, indicating that the amount of gait data is large
Data distribution. Fig. 12 shows the data distribution of AddBiomechanics (Werling et al., 2024). As shown, over 75% of the data are either walking, running, or standing, which are extremely limited. Though ImDyS is empowered with the diverse ImDy, it still requires data to learn the mapping between simulated torques and real torques for non-gait data. These result in ImDyS’ poor performance when processing non-gait data. Further mitigating the limited data issue for non-gait motions would be a meaningful goal to pursue.

### v2

@@ 1 Introduction
The rapid progress in human motion capture based on computer vision has made an enormous amount of human motion data available to the research community (Mahmood et al., 2019; Mandery et al., 2016). The accumulation of human motion manages to push motion understanding forward in various tasks, including behavior understanding (Punnakkal et al., 2021; Shahroudy et al., 2016) and character animation (Guo et al., 2022; Tevet et al., 2023; Liu et al., 2025a; b). However, given the vision-based nature, most current efforts focus only on visible kinematics information. The invisible factors, especially the dynamic factors, which could carry deeper insights into the underlying production mechanism of human motion, are typically overlooked, such as driven torques and ground reaction forces. This limits the current motion understanding algorithms from wider applications to domains where physical constraints must be seriously considered, such as robotics (Figueredo et al., 2020; Teramae et al., 2017), healthcare (Yao et al., 2018), and sports training  (Caruntu & Moreno, 2019). To alleviate this, we focus on identifying the driven torques and ground reaction forces for human motion from pure kinematics MoCap data, known as human inverse dynamics (ID).
Human inverse dynamics, as a basic step toward physical motion modeling, has been extensively discussed by the biomechanics community for applications like gait analysis. A fundamental obstacle is that it could not be measured non-intrusively. Therefore, computationally expensive optimization-based methods are widely adopted and mature software is developed (Delp et al., 2007; Damsgaard et al., 2006; Werling et al., 2021). However, accurately measured ground reaction forces are required to ensure a determinate solution, which could be expensive and applicable only in restricted laboratory settings. Also, the optimization process could be sensitive to small disturbances in either motion capture noises or subject variances. These make it hard to scale up for wider applications to general motion. Given the success achieved by data-driven methods in CV and NLP, deep-learning-based methods are proposed (Zell & Rosenhahn, 2015; Zell et al., 2017; Lv et al., 2016), aiming at scalable human inverse dynamics with only kinematic observations as inputs.
@@
Given this, we project our sights on the recent progress of Imitation Learning (IL) (Luo et al., 2021; 2023), which replicates recorded human motion through fully simulated humanoids with physical control signals, namely, joint torques. A key insight is that with the goal of kinematics phenomenon imitation, IL might also implicitly imitate the dynamics production mechanism, known as ID. However, IL is not directly applicable to ID. Despite the visual resemblances between the recorded and simulated motion, kinematic errors still exist. These errors could be neglected for kinematic analyses, however, for dynamic analysis, they could be amplified drastically (Uchida & Seth, 2022). Moreover, existing successful IL algorithms are typically based on joint-actuated SMPL (Loper et al., 2015) avatars, whose physical properties and topology differ from real humans. To this end, extracting ID knowledge from IL becomes critical. Here, we adopt the state-of-the-art motion IL algorithm (Luo et al., 2023) and physics simulator (Makoviychuk et al., 2021) to imitate recorded motions, extracting the observed kinematic states, joint torques, and the ground reaction forces, resulting in a large-scale human inverse dynamics database named Imitated Dynamics (ImDy) with more than 150-hour human motion. There are two major merits of ImDy. First, it is scalable. Multiple samples could be concurrently collected in the simulator without expensive laboratory setups, extending the border of ID data acquisition. As shown in Fig. 1, we could even pair some rather complex motions with ID data, which is hard to achieve in laboratories. Second, it is holistic. Beyond the ground reaction force and ID typically recorded in laboratories for previous efforts (Zell et al., 2020; Mourot et al., 2022; Han et al., 2023), the physics simulator enables us to access the GRFs and joint torques of all human body segments, as shown in Fig. 1.
@@
We demonstrate the efficacy of ImDyS through a wide span of experiments. First, we evaluate our method on ImDy for a basic performance illustration with simulated ImDy. Then ImDyS is evaluated on GroundLink (Han et al., 2023), which contains real-world ground reaction force. Furthermore, we demonstrate the efficacy of ImDy on the recent real-world human dynamics dataset AddBiomechanics (Werling et al., 2025).
@@ 2 Background
Conventional Inverse Dynamics. Inverse dynamics, known as inferring forces/moments from kinematic observations, have been discussed for long in the biomechanics community. In this literature, it is formulated as an optimization problem: given a representative model of a subject, the joint kinematics over time w.r.t. the subject model, and the external forces, find the driving torques that produce the motion (Uchida & Delp, 2021). The Newtonian dynamic equations are involved as
@@ 3.2 Imitated Data Acquisition
AddBiomechanics (Werling et al., 2025) 273 57.6 GRF* & Torques Rajagopal et al. (2016) Real
@@
We adopt AMASS (Mahmood et al., 2019) and KIT (Krebs et al., 2021) as two major data sources. Sequences involving humans interacting with objects other than the ground are excluded, resulting in over 50 hours of motion. All the sequences are re-sampled to 30FPS, with the z-axis as the gravity axis. Then, the sequences are imitated three times by the two primitive policies and the multiplicative policy with a simulation frequency of 60Hz, resulting in over 150 hours of human motion data with dynamics. States including $q,\dot{q},\beta$ are recorded in synchronous with the torque $\tau$, all restored in the format of SMPL (Loper et al., 2015) if possible. Moreover, GRFs for the whole body are also recorded, resulting in ImDy, a large-scale human motion dynamics dataset.
@@ 4.1 Formulation
Motion States $s$ could be represented by either SMPL parameters, joint angles, joint coordinates, or marker coordinates. However, due to the topology divergence, the conversion among SMPL parameters, joint angles, and joint coordinates is non-trivial with limited performances. To guarantee that ImDyS could be seamlessly adopted to both ImDy and real-world biomechanics data, we adopt marker coordinates as motion state representation for ImDyS. The state $s^{t}=(m^{t},\dot{m}^{t})$ is composed of marker coordinates $m^{t}$ and finite-differentiated velocities $\dot{m}^{t}$ at timestamp $t$, which are easy to obtain for both ImDy and AddBiomechanics (Werling et al., 2025). Two temporal windows before and after the transition with a length of $w$ are included for contextual information. Notice that human physical properties like height and weight could also be implicitly represented by the markers. The states are canonicalized w.r.t. the heading direction of $s^{t}$.
@@
Ground Reaction Forces $\lambda$. Different from previous efforts (Mourot et al., 2022; Han et al., 2023; Werling et al., 2025) with foot GRFs only, we predict full-body GRF $\lambda\in\mathbb{R}^{J\times 3}$ as in Fig. 1.
@@ 5.4 Evaluation on AddBiomechanics
Metrics. AddBiomechanics (Werling et al., 2025) is recently proposed with over 50 hours of human dynamics data from 273 subjects. We adopt the armless part of this dataset. We follow the train/test split in Addbiomechanics and report mPJE for the joint torque normalized by body weight.
@@ 7 Conclusion
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China under Grants No.62302296, 62306175, and 62472282.
@@ Appendix B AddBiomechanics Results Analysis
Figure 12: Data distribution of AddBiomechanics and ImDy. Among all activities, walking and running account for over 75% of AddBiomechanics. In comparison, according to the annotations from BABEL (Punnakkal et al., 2021), ImDyS is less imbalanced with better diversity.
Data distribution. Fig. 12 shows the data distribution of AddBiomechanics (Werling et al., 2025). As shown, over 75% of the data are either walking, running, or standing, which are extremely limited. Though ImDyS is empowered with the diverse ImDy as shown in Fig. 12, it still requires data to learn the mapping between simulated torques and real torques for non-gait data. These result in ImDyS’ poor performance when processing non-gait data. Further mitigating the limited data issue for non-gait motions would be a meaningful goal to pursue.
@@
Appendix C Sim2Real Analysis
Figure 15: Knee torque magnitude visualization of ImDyS and ImDyS w/o Sim2Real fine-tuning on AddBiomechanics. ImDyS w/o Sim2Real produces larger magnitudes and over-active torques w/o Sim2Real fine-tuning as circled in red.
We further analyze the Sim2Real effect of ImDy(S) via Fig. 15. An interesting question is the performance of ImDyS without any fine-tuning on AddBiomechanics. Though this could be inapplicable for most joints due to the human model definition discrepancy between Rajagopal’s model in AddBiomchanics and SMPL in ImDy, the knee joints in the two models could roughly correspond to each other. Therefore, we visualize the knee torque magnitudes of ImDyS and ImDyS w/o Sim2Real finetuning on AddBiomechanics in Fig. 15. Even without fine-tuning, ImDyS could reproduce the trends of knee torque magnitudes. However, artifacts could also be observed in two aspects. First, ImDyS w/o Sim2Real tends to produce much larger torques. Second, ImDyS w/o Sim2Real could be over-active compared to real humans and ImDyS like in the red circles. The reason could be two-fold. First, the simulation parameters used by ImDy, like mass and inertia, are different from real humans. Second, though the knee joints could roughly correspond, the knee in SMPL has more DoFs than Rajagopal’s model, which might require larger torques to produce similar motions. With the simple Sim2Real fine-tuning of ImDyS, the issues could be alleviated. Further exploration for better Sim2Real performance would be meaningful future work.
Appendix D Details on Data Flow
Figure 16: Details of $L_{prior}$ and $L_{FD}$.
Details of the adopted $L_{prior}$ and $L_{FD}$ are illustrated in Fig. 16.
For $L_{prior}$, the input motion state is treated as the positive case, and we generate corresponding negative cases by either temporal permutation or adding random noises. The samples are fed to the encoder, and the prior discriminator predicts whether the sample is positive.
For $L_{FD}$, we first feed ImDyS with motion state $s^{t-w:t+w+1}$, obtaining $\tau,\lambda$. Then, $\tau,\lambda,s^{t-w:t}$ are fed into the FD model, outputing $\hat{s}^{t+1}$. The FD loss is computed as $L_{FD}=|s^{t+1}-\hat{s}^{t+1}|$.
Appendix E Analysis on FD Model
We report the marker RMSE of the FD model on the AddBiomechanics test set as Tab. 8. ImDyS noticeably outperforms the baseline trained on AddBiomechanics only, indicating the importance of ImDy pre-training. Moreover, ImDyS is competitive even compared to the differentiable simulator Nimble.
Table 8: Extended results on the FD model on Addbiomechanics.
Methods Baseline ImDyS Nimble
RMSE 0.0302 0.0194 0.0186
Appendix F Comparison with Original PHC
Due to the inaccessible torques, we did not include the original PHC as a baseline for ImDy. However, it is noticeable that the original PHC can also conduct GRF prediction. To this end, we also evaluate the original PHC on GroundLink. It provides a left-foot mPJE of 1.559 and a right-foot mPJE of 3.518, which are comparable to re-trained PHC and worse than our proposed ImDyS. Some visualizations are included in Fig. 17. Even without the naive PD controller, the original PHC could suffer from jittering predictions, which could result from the non-perfect contact simulation. In contrast, ImDyS could produce smoother predictions with higher precision.
Figure 17: Original PHC on GroundLink.